OpenStudy (anonymous):

what is the equation of the line perpendicular to the line y=4/9x-2 that passes through the point (4,3)?

OpenStudy (shamim):

U hv an equation of a line is y=(4/9)x-2 Right?

OpenStudy (shamim):

Standard equation of a line is y=mx+c

OpenStudy (shamim):

m=slope

OpenStudy (shamim):

So slope of ur given line=m1=4/9 Right?

OpenStudy (shamim):

Let the slope of ur perpendicular line =m2

OpenStudy (shamim):

So we can write m1*m2=-1

OpenStudy (shamim):

(4/9)*m2=-1 m2=-9/4

OpenStudy (shamim):

Use y-y1=m2(x-x1)